France and Germany each produce both wine and beer. Their production possibilities are: France - 200 bottles of wine or 400 bottles of beer, and Germany 500 bottles of wine or 400 bottles of beer.

A. Which country has an absolute advantage in wine production?

B. Which country has an absolute advantage in beer production?

C. Which country has a comparative advantage in wine production?

D. Which country has a comparative advantage in beer production?

E. Propose a trade agreement between the two countries that each would likely accept rather than remain self-sufficient and provide justification for your answer.
Answer #1

A. Germany has an absolute advantage in producing wine
B. No country has an absolute advantage in producing beer.

C. France's opportunity cost of producing wine = 400/200= 2 beer

Germany's opportunity cost of producing wine= 400/500= 0.8 beer

Germany has a comparative advantage in wine

D. France has a comparative advantage in beer.

E. Terms of trade should be in between 0.8 beer and 2 beer for 1 bottle of wine.

Because this will make both countries better off when they specialize and trade